Suppose that the middle 95% of average scores in the NBA per player per game fall between 8.31 and 29.15. Give an approximate estimate of the standard deviation of the number of the points scored. Assume the points scored has a normal distribution.

Options:

1)

2.605

2)

10.42

3)

-5.21

4)

5.21

5)

18.73

Suppose that the distribution of income in a certain tax bracket is approximately normal with a mean of $46,188.21 and a standard deviation of $1,438.254. Approximately 56.46% of households had an income greater than what dollar amount?

Options:

1)

45,954.29

2)

We do not have enough information to calculate the value.

3)

382,626.2

4)

290,249.8

5)

46,422.13
Answer #1

We are allowed to do only 1 question at a time. Post again for other answers.

1)

fall between 8.31 and 29.15

Mean - 2 * SD = 8.31

Mean + 2 * SD = 29.15

4 * SD = 20.84

SD = 5.21